What happens to contribution margin per unit, contribution margin ratio, contribution margin in total, and net operating income when activity levels increase?
Answer #1

contribution margin per unit is calculated as selling price per unit minus variable cost per unit.there will be no change arises when the activity level increases.consider an example ABC co sells a product at selling price $20 per unit and variable cost per unit is $12.the company expencts to sell 10000 units.the contribution margin per unit is $8 ($20-$12).if the company increases its selling units to 12000 units the contribution margin per unit won't change.

contribution margin ratio is calculated as total contribution is divided by total sales.the contribution margin ratio will not change for the increase in activity level.consider an example ABC co sells a product at selling price $20 per unit and variable cost per unit is $12.the company expencts to sell 10000 units.

contribution margin ratio=total contribution/tota lsales

total contribution=$8*10000=$80000

total sales=$20*10000=$200000

contribution margin ratio=$80000/$200000=0.4

.if the company increases its selling units to 12000 units the contribution margin ratio

total contribution=$8*12000=$96000

total sales=$240000

contribution margin ratio=$96000/$240000=0.4

contribution margin in total is calaculated as multiplying the contribution per units with sales units.contribution margin in total will increase with increase in activity level.consider an example ABC co sells a product at selling price $20 per unit and variable cost per unit is $12.the company expencts to sell 10000 units.the contribution margin per unit is $8 ($20-$12).the contribution margin intotal is $80000 ($8*10000) and if the company increases its selling units to 12000 units the contribution margin in total is $96000 ($8*12000).

net opearting income is calculated by reducing the fixed cost from contribution margin in total.the net opreating income will inreases with the increase in activity level.consider an example ABC co sells a product at selling price $20 per unit and variable cost per unit is $12.the company expencts to sell 10000 units.the contribution margin per unit is $8 ($20-$12).the contribution margin intotal is $80000 ($8*10000) and if the company have fixed cost of $50000, then the net operating income will be $30000 ($80000-$50000).if the company increases its selling units to 12000 units the contribution margin in total is $96000 ($8*12000).the company have the same fixed cost of $50000 beacuse the fixed cost won't change with the change in activity level.the net operating income will be $46000 ($96000-$50000).